# A Secure Estimator with Gaussian Bernoulli Mixture Model

### v1

@@ Abstract
The implementation of cyber-physical systems in real-world applications is challenged by safety requirements in the presence of sensor threats. Most cyber-physical systems, in particular the vulnerable multi-sensor systems, struggle to detect the attack in observation signals. In this paper, we tackle this issue by proposing a Gaussian-Bernoulli Secure (GBS) estimator, which effectively transforms the assessment of sensor status into an optimal estimation problem concerning the system state and observation indicators. It encompasses two theoretical sub-problems: sequential state estimation with partial observations and estimation updates with disordered new observations. Within the framework of Kalman filter, we derive closed-form solutions for these two issues. However, due to their computational inefficiency, we propose the iterative approach employing proximal gradient descent to accelerate the estimation update. We conduct comprehensive experiments from three perspectives: computational efficiency, detection and estimation performance, and characterization of observation error. Our GBS estimator shows the improvements compared to other methods.
Index Terms:
@@
I Introduction
Cyber-Physical Systems (CPSs) are integrative frameworks that combine computing and communication capabilities to monitor or control entities within the physical world [1][2]. Over the past two decades, significant advancements in communication and computing technologies have catalyzed the wide applications for CPSs [3]. The advancement in wireless communication technologies has enabled CPSs to transmit real-time signals across expansive physical networks [4]. Furthermore, armed with advanced computational resources, CPS can optimize control strategies for complex systems online in real-time  [5]. This technological evolution broadens the scope of CPSs, impacting various sectors such as medical devices [6], smart grids [7], and autonomous vehicles [8].
@@ I Introduction
Against various sensory attacks, two types of detection methods are widely studied. One is the model-based defense mechanisms, e.g., control theory, game theory and cryptography. Detectors based on linear estimation theory, which analyze the Kalman innovation sequence, are widely applied in CPS, especially windows $\chi^{2}$ detector [15], CUSUM detector [16] and MEWMA detector [17]. The framework of game theory also enables the design of the optimal defense policies. For instance, Basar et al. [5] considered the system balance between resilience and robustness, Li et al. [18] obtained the optimal DoS defense strategy by solving the associated Bellman equations, and Meira-Góes et al. [19] improved the efficiency of synthesized robust supervisors against sensor deception attacks. Additionally, with the idea of cryptography, Mo et al. [20] proposed the watermarking approach to authenticate the correct operation of a control system, while Kogiso et al. [21] incorporated the public key encryption scheme into the design of remote controller or observers. Another strategy involves using learning-based methods to identify CPS attacks. The graph neural network (GNN) approach proposed in [22] provided a solution for interpretation and localization of anomalies in secure water treatment, overcoming the challenges of high dimensionality and complexity. Besides, LSTM-RNNs were applied in [23] for learning the temporal behaviour of the data in CPSs and identifying sensor anomalies. Moreover, Wang et al. [24] utilized Deep Belief Network (DBF) to train an interval state predictor for pattern recognition in sensor security.
@@
Apart from detecting attacks, designing resilient state estimators played a critical role in mitigating sensor attacks. Mo et al. [25] studied the case of multiple measurements subject to sensor attacks, designing an optimal estimator by solving a minimax optimization problem. In addition, regarding secure estimation problem in noiseless systems, Chang et al. [26] formulated it as the classical error correction problem and provided the sufficient conditions for restoring the real states. Furthermore, Wang et al. [27] considered attacked sensor signals as general outliers, obtaining resilient estimates by identifying and eliminating these outliers. Previous detecting methods have two limitations. First, attack detectors and resilient state estimators are interdependently designed and the connection between them are overlooked. It is challenging to ensure the estimation performance while simultaneously detecting attacks. Second, occasional outliers and persistent malicious attacks are difficult to distinguish, as their Kalman innovation sequences may overlap. If we can separate these outliers from observations, the binary detection problem will become more accurate, thereby reducing the false alarm rate.
@@
    We introduce, for the first time, the Gaussian-Bernoulli mixture model to model the observation model with potential attack risks. Under the mixture model, we propose a novel framework to integrate the attack detection and resilient estimation into a dual-variable optimization problem.
@@
    We propose an innovative iterative method that addresses the challenge of non-sequential observation changes. By using the previous estimation as the initial point and employing proximal gradient descent, it rapidly converges to the optimal estimation with updated observations.
@@
    We test the performance of the GBS estimator under various types and intensities of attacks, and observe a significant improvement compared to previous algorithms.
@@
II Bayes’ theorem in HMMs
This section introduces Bayesian estimation as tools for analyzing stochastic systems and estimating the states of a hidden Markov model (HMM). Furthermore, we present the steps of Kalman filter and RTS smoother, highlighting their optimality within the theoretical framework.
II-A Bayesian estimation
Bayesian estimation is a statistical inference method for updating the beliefs in presence of new evidence. This approach is particularly useful where data is limited or uncertainty is high.
At the heart of Bayesian estimation is Bayes’ theorem, which describes the relationship between conditional probabilities. It applies the concepts of probability to the random variables and their relations, including the prior probability $p(x)$, likelihood function $p(y|x)$ and posterior probability $p(x|y)$. The theorem can be expressed as:
$p(x|y)=\frac{p(y|x)\cdot p(x)}{p(y)}$
where $p(x)$ represents current knowledge before observation, $p(y|x)$ describes the known statistical model of observation given the state, and $p(x|y)$ updates probability of the state after observation.
@@ II-A Bayesian estimation
II-A1 MMSE
In the Bayesian framework, the goal of MMSE estimation is to minimize the mean squared error between the estimated values and the true values. Mathematically, the MMSE estimator can be expressed as:
$\hat{x}^{\text{MMSE}}=\arg\min_{\hat{x}}\mathbf{E}[\|x-\hat{x}\|^{2}|y]$
Further, when the posterior distribution is symmetric without the heavy tail, the MMSE estimator is equal to the expectation under certain conditions.
[…]
The MMSE estimator focuses on minimizing the overall error, providing an estimate that performs best on average across all possible values.
II-A2 MAP
The MAP estimator aims to find the state value that maximizes the posterior probability. It means that given the observed data, we seek the state value most likely to produce these data. Mathematically, the MAP estimator can be expressed as:
$\hat{x}^{\text{MAP}}=\arg\max_{\hat{x}}p(\hat{x}|y)=\arg\max_{\hat{x}}\frac{p(y|\hat{x})\cdot p(\hat{x})}{p(y)}$
MAP estimator aims to determine the most probable parameter value, providing a more convicing estimate when the random variables are discrete.
II-B Kalman filter and RTS smoother
@@
First, both state transitions and observation processes are assumed to be linear and accompanied by Gaussian noise. They can be expressed as:
$x_{t+1}\sim\mathcal{N}(A{x}_{t},Q)$
$y_{t}\sim\mathcal{N}(C{x}_{t},R)$
where $Q>0$ and $R>0$ denotes the process noise and observation noise. The probability density function of $\mathcal{N}(\mu,\Sigma)$ is given by the following formula
$p(x)\propto\exp\left(-\frac{1}{2}\|x-\mu\|^{2}_{\Sigma^{-1}}\right).$
@@
$\hat{x}^{\text{LMMSE}}_{t},F=\begin{array}[]{clll}\displaystyle\arg&\min_{\hat{x}_{t},F}E[\|x_{t}-\hat{x}_{t}\|^{2}]\\ \textrm{s.t.}&\hat{x}_{t}=A\hat{x}_{t|t}^{*}+F(\hat{x}^{\text{LMMSE}}_{t+1}-A\hat{x}_{t|t}^{*})\end{array}$
@@
Moreover, LMMSE and MAP have the same result in linear Gaussian model. The closed-form solution of optimal estimate can be calculated by Kalman filter and Rauch-Tung-Striebel (RTS) smoother, respectively.
@@
Kalman filter
$\hat{x}_{t+1|t}$ $=A\hat{x}_{t|t}$
$P_{t+1|t}$ $=AP_{t|t}A^{T}+Q$
$K_{t+1}$ $=P_{t+1|t}C^{T}(CP_{t+1|t}C^{T}+R)^{-1}$
$r_{t+1}$ $=y_{t+1}-C\hat{x}_{t+1|t}$
$\hat{x}_{t+1|t+1}$ $=\hat{x}_{t+1|t}+K_{t+1}r_{t+1}$
$P_{t+1|t+1}$ $=(I-K_{t+1}C)P_{t+1|t}$
RTS smoother
$F_{t}$ $=P_{t|t}A^{T}P_{t+1|t}^{-1}$
$\hat{x}_{t|t+1}$ $=\hat{x}_{t|t}+F_{t}(\hat{x}_{t+1|t+1}-A\hat{x}_{t|t})$
$P_{t|t+1}$ $=P_{t|t}+F_{t}(P_{t+1|t+1}-P_{t+1|t})F_{t}^{T}$
III problem formulation
@@
III-A System model
@@
where the system state $x_{t}\in\mathbf{R}^{n}$, the measurement $y_{t}\in\mathbf{R}^{m}$ and the pair $(A,C)$ is observable. We assume the process noise $\{w_{t}\}$ satisfies i.i.d Gaussian distribution $\mathcal{N}(0,Q)$, and is not uncorrelated with measurement noise $\{\bar{v}_{t}\}$, which is non-Gaussian, especially when under attack.
@@
The challenge is how to model $\Delta v_{t}$. There are two intuitions for modeling $\Delta v_{t}$. One is that the fitting error between $v_{t}$ and the real sample data is relatively uniform across different values. Another motivation is that it is difficult to accurately characterize the distribution of large error events, but it is straightforward to use a Bernoulli distribution to describe the likelihood of one of large error events occurring.
@@
where $\{v_{t}\}$ has Gaussian ditribution $\mathcal{N}(0,R)$, $\{p_{t}\}$ has Bernoulli distribution $\mathcal{B}(N,\beta)$, and $\{\delta_{t}\}$ has uniform distribution with unknown probability density. Random variables $\{p_{t}\},\{\delta_{t}\},\{v_{t}\}$ are uncorrelated with each other.
This mixture model can be interpreted as conforming to a Gaussian distribution in most cases with occasional unknown distribution. We name the $\{p_{t}\}$ as observation indicators.
III-B Problem description
Our goal is to design a secure detector to determine whether the observation sequence of fixed length $N$ is reliable. The secure detector operates in parallel with the estimator that updates the state estimation. In other words, at a certain moment, the secure detector can use the observations before and after to determine whether it is reliable.
@@ III-B Problem description
Figure 1: Illustration of the detection problem under sensor attacks.
For each sensor, the evaluation of the observation sequence relies on the inference of each observation. Under Assumption 1, $y_{t}$ is considered normal when $p_{t}=0$, as the observation noise conforms to a Gaussian distribution; whereas when $p_{t}=1$, the observation noise is with unknown distribution, which is considered abnormal. Based on the $0-1$ sequence of observation indicators $p_{0:N}$, we can classify the observation sequence into three status with threshold $\tau$: in normal case, with general outliers, or under malicious sensor attacks.
$\{p_{0},\cdots,p_{N}\}$ $\sum_{i=0}^{N}p_{t}$ the status of observation sequence
$0,0,0,\cdots,0$ 0 in normal case
$0,1,0,\cdots,0$dots $\leq\tau$ with general outliers
$\dots$ $1,1,1,\cdots,1$ $>\tau$ under sensor attacks
@@
The problems of obtaining $\{\hat{p}_{t}\}$ and $\{\hat{x}_{t}\}$ are coupled. A reliable detector relies on accurate state estimator, while at the same time, effective elimination of anomalies contributes to better estimation performance. Consequently, we formulate it to a dual-variable optimization problem
@@
It is a Mixed-Integer Programming (MIP) problem with a mix of continuous variables ($\hat{x}_{0:N}$) and Boolean variables ($\hat{p}_{0:N}$). Solving such problem is challenging due to the non-convex nature of the objective function and the combinatorial complexity introduced by integer variables.
@@
IV Direct approach for state estimation
In this section, we solve a special case of Problem 1. When the observation indicators are fixed, what is the closed-form solution of equation (7)? Furthermore, when observation indicators changes, how should the solution of equation (7) be updated? These are the two key steps of our algorithm.
When $\hat{p}_{0:N}$ is fixed, let us define the selected index set $\mathcal{O}=\{i|\hat{p}_{i}=0\}$. Problem 1 degenerates into the following problem.
@@ IV Direct approach for state estimation
As mentioned in Section II-B, the solutions for MAP and LMMSE estimation in linear Gaussian systems are the same. Therefore, in this section, we present a direct solution to equation (8) from the perspective of LMMSE.
IV-A The estimation of state sequence with partial observations
@@
Define the column vectors $X=col\{x_{0},\cdots,x_{N}\},Y=col\{y_{0},\cdots,y_{N}\}$ and $\hat{X}_{\mathcal{O}}=col\{\hat{x}_{0|\mathcal{O}},\cdots,\hat{x}_{N|\mathcal{O}}\}$. Problem 3 describes our optimization objectives and conditions.
@@
$\begin{array}[]{clll}\displaystyle\arg\min_{\hat{X}_{\mathcal{O}},K_{\mathcal{O}}}&J_{\mathcal{O}}=E[\|X-\hat{X}_{\mathcal{O}}\|^{2}]\\ \textrm{s.t.}&\hat{X}_{\mathcal{O}}=K_{\mathcal{O}}\cdot Y\\ \end{array}$ (9)
where each element in $X$ and $Y$ satisfies
$x_{t+1}$ $\sim\mathcal{N}(Ax_{t},Q),\hskip 0.5ptt=0,\cdots,N,$
@@
This problem is equivalent to finding the optimal estimation for each state individually. Based on the theory of linear estimation, the information from $y_{\mathcal{O}}$ can be well utilized by the Kalman filtering process and RTS smoother. The update of state estimation on the sequence follows this rule: deliver the past information forwards through prediction, fuse the sensor information through measurement update, and supplement future information backwards through the smoother. It is worth mentioning that for those states missing corresponding observations, we can skip the step of measurement update.
@@
    Calculate $\{K_{i}\}$,$\{F_{i}\}$ according to the Kalman filter and RTS smoother process. For those intermittent observations that $\{y_{j}|j\notin\mathcal{O}\}$, set the Kalman gain to zero $\{K_{j}=0|j\notin\mathcal{O}\}$.
@@
The proof of Lemma 1 is in the appendix.A-A. Moreover, Lemma 1 provides the explicit linear relationship between estimated states $\hat{x}_{0:N|\mathcal{O}}$ and the observations $y_{\mathcal{O}}$. For those $\{y_{i}|i\notin\mathcal{O}\}$, when multiplied by the matrix $K_{\mathcal{O}}$, the corresponding weights are zeros, which clearly indicates that they are unrelated to the estimate.
IV-B Estimation update for the disordered new observation

### v2

# Secure State Estimation of Cyber-Physical Systems via Gaussian Bernoulli Mixture Model

@@ Abstract
The implementation of cyber-physical systems in real-world applications is challenged by safety requirements in the presence of sensor threats. Most cyber-physical systems, especially multi-sensor systems, struggle to detect sensor attacks when the attack model is unknown. In this paper, we tackle this issue by proposing a Gaussian-Bernoulli Secure (GBS) estimator, which transforms the detection problem into an optimal estimation problem concerning the system state and observation indicators. It encompasses two theoretical sub-problems: sequential state estimation with partial observations and estimation updates with disordered new observations. Within the framework of Kalman filter, we derive closed-form solutions for these two problems. However, due to their computational inefficiency, we propose the iterative approach employing proximal gradient descent to update the estimation in less time. Finally, we conduct experiments from three perspectives: computational efficiency, detection performance, and estimation error. Our GBS estimator demonstrates significant improvements over other methods.
keywords:
@@
1 Introduction
Cyber-Physical Systems (CPSs) are integrative frameworks that combine computing and communication capabilities to monitor or control entities within the physical world [1, 2, 3]. Over the past two decades, significant advancements in communication and computing technologies have catalyzed the wide applications for CPSs [4]. The advancement in wireless communication technologies has enabled CPSs to transmit real-time signals across expansive physical networks [5]. Furthermore, armed with advanced computational resources, CPS can optimize control strategies for complex systems online in real-time [6]. This technological evolution broadens the scope of CPSs, impacting various sectors such as robot manipulators [7], smart grids [8], and autonomous vehicles [9], to name just a few.
@@ 1 Introduction
Against various sensory attacks, two types of detection methods are widely studied. One is the model-based defense mechanisms, e.g., control theory, game theory and cryptography. Detectors based on linear estimation theory, which analyze the Kalman innovation sequence, are widely applied in CPS, especially windows $\chi^{2}$ detector [16], interval anomaly detector [17] and LASSO-based detector [18], to name just a few. The framework of game theory also enables the design of the optimal defense policies. For instance, Basar et al. [19] considered the system balance between resilience and robustness, Li et al. [20] obtained the optimal DoS defense strategy by solving the associated Bellman equations, and Meira-Góes et al. [21] improved the efficiency of synthesized robust supervisors against sensor deception attacks. Fu et al proposed a distributed secure filtering method against eavesdropping attacks in SINR-based sensor networks [22]. Additionally, with the idea of cryptography, Mo et al. [23] proposed the watermarking approach to authenticate the correct operation of a control system, while Kogiso et al. [24] incorporated the public key encryption scheme into the design of remote controller or observers. The encryption-based secure method is further discussed in [25, 26, 27]. Another strategy involves using data-driven methods to identify CPS attacks. The graph neural network (GNN) approach proposed in [28] provided a solution for interpretation and localization of anomalies in secure water treatment, overcoming the challenges of high dimensionality and complexity. Besides, LSTM-RNNs were applied in [29] for learning the temporal behaviour of the data in CPSs and identifying sensor anomalies. Moreover, Wang et al. [30] utilized Deep Belief Network (DBF) to train an interval state predictor for pattern recognition in sensor security. Reinforcement learning methods have also been utilized in online detection and mitigation of CPS attacks [31].
@@
Apart from detecting attacks, designing resilient state estimators played a critical role in mitigating sensor attacks. Mo et al. [32] studied the case of multiple measurements subject to sensor attacks, designing an optimal estimator by solving a minimax optimization problem. In addition, regarding secure estimation problem in noiseless systems, Chang et al. [33] formulated it as the classical error correction problem and provided the sufficient conditions for restoring the real states. Furthermore, Wang et al. [34] considered attacked sensor signals as general outliers, obtaining resilient estimates by identifying and eliminating these outliers. The resilient method is also studied in the consensus of multi-agent systems [35] and distributed estimation [36]. Previous detecting methods have two limitations. First, attack detectors and resilient state estimators are interdependently designed and the connection between them are overlooked. It is challenging to ensure the estimation performance while simultaneously detecting attacks. Second, occasional outliers and malicious attacks are difficult to distinguish, as their Kalman innovation sequences may overlap. If we can separate these outliers from observations, the binary detection problem will become more accurate, thereby reducing the false alarm rate.
@@
    In the absence of knowledge about the attack model, we introduce the Gaussian-Bernoulli mixture model to represent the observation model. Under this mixture model, we propose a novel framework to integrate the attack detection and resilient estimation into an optimal estimation problem with dual variables.
@@
    We propose an innovative iterative method that addresses the challenge of non-sequential observation changes. By using the previous estimation as the initial point, our algorithm utilizes proximal gradient descent to accelerate convergence to the optimal estimation under the new observations.
@@
    We test the performance of the GBS estimator without any attacker information, across different types and intensities of attacks. The simulations show that it significantly outperforms previous methods in both detection and estimation performance.
@@
2 Bayesian Methods in HMMs
This section introduces Bayesian estimation as tools for analyzing stochastic systems and estimating the states of a hidden Markov model (HMM). Furthermore, we present the steps of Kalman filter and RTS smoother, highlighting their optimality within our framework.
2.1 Bayesian Estimation
Bayesian estimation is a statistical inference method for updating the beliefs in presence of new evidence. This approach is particularly valuable when data is limited or uncertainty is high.
At the core of Bayesian estimation is Bayes’ theorem, which describes the relationship between conditional probabilities. The theorem can be expressed as:
$p(x|y)=\frac{p(y|x)\cdot p(x)}{p(y)},$
where the prior probability $p(x)$ represents current knowledge before observation, the likelihood function $p(y|x)$ describes the known statistical model of observation given the state, and the posterior probability $p(x|y)$ updates probability of the state after observation.
2.2 MMSE and MAP
@@
The goal of MMSE estimation is to minimize the mean squared error between the estimated values and the true values. Mathematically, the MMSE estimator can be expressed as:
$\hat{x}^{\text{MMSE}}=\arg\min_{\hat{x}}\mathbf{E}[\|x-\hat{x}\|^{2}|y].$
Further, when the posterior distribution is symmetric without the heavy tail, the MMSE estimator is equal to the expectation under certain conditions,
$\hat{x}^{\text{MMSE}}=\mathbf{E}[x|y]=\int x\cdot\frac{p(y|x)\cdot p(x)}{p(y)}dx.$
In contrast to the MMSE estimator, the MAP estimator aims to find the state value that maximizes the posterior probability. It means that given the observed data, we seek the state value most likely to produce these data. Mathematically, the MAP estimator can be expressed as:
$\hat{x}^{\text{MAP}}=\arg\max_{\hat{x}}p(\hat{x}|y)=\arg\max_{\hat{x}}\frac{p(y|\hat{x})\cdot p(\hat{x})}{p(y)}.$
2.3 Kalman filter and RTS smoother
@@
First, both state transitions and observation processes are assumed to be linear and accompanied by Gaussian noise. They can be expressed as follows,
$x_{t+1}\sim\mathcal{N}(A{x}_{t},Q),$
$y_{t}\sim\mathcal{N}(C{x}_{t},R),$
where $Q>0$ and $R>0$ denotes the process noise and the observation noise.
@@
$\hat{x}^{\text{LMMSE}}_{t},F=\begin{array}[]{clll}\displaystyle\arg&\min_{\hat{x}_{t},F}E[\|x_{t}-\hat{x}_{t}\|^{2}]\\ \textrm{s.t.}&\hat{x}_{t}=\hat{x}_{t|t}^{*}+F(\hat{x}^{\text{LMMSE}}_{t+1}-A\hat{x}_{t|t}^{*})\end{array}$
@@
Moreover, LMMSE and MAP have the same result in linear Gaussian model. The closed-form solution of optimal estimate can be calculated by Kalman filter and Rauch-Tung-Striebel (RTS) smoother.
@@
3 Problem Formulation
@@
3.1 Cyber Physical System Model
@@
where $x_{t}\in\mathbf{R}^{n}$ is the state, $y_{t}\in\mathbf{R}^{m}$ the measurement, and the pair $(A,C)$ is observable. We assume the process noise $\{w_{t}\}$ satisfies i.i.d Gaussian distribution $\mathcal{N}(0,Q)$, and is not uncorrelated with measurement noise $\{\bar{v}_{t}\}$, which is non-Gaussian, especially when under attack.
@@
The challenge is how to model $\Delta v_{t}$. There are two intuitions for modeling $\Delta v_{t}$. On the one hand, since no prior information about external attacks is available, it is reasonable to assume that the attack signals follow a non-informative distribution over a finite interval. On the other hand, although the exact distribution of outliers is difficult to characterize, the occurrence of abnormal events can be modeled using a Bernoulli distribution.
@@
where $\{v_{t}\}$ has Gaussian ditribution $\mathcal{N}(0,R)$, $\{p_{t}\}$ has Bernoulli distribution $\mathcal{B}(N,\beta)$, and $\{\delta_{t}\}$ has uniform distribution with unknown probability density. Random variables $\{p_{t}\},\{\delta_{t}\},\{v_{t}\}$ are uncorrelated.
This mixture model can be interpreted as conforming to a Gaussian distribution with occasional unknown distribution. We regard $\{p_{t}\}$ as observation indicators.
3.2 Problem of Interest
Our goal is to design a secure detector to determine whether the sequential observations of fixed length $N$ are reliable. The secure detector operates in parallel with the estimator that updates the state estimation. In other words, at a certain moment, the secure detector can use the observations before and after to determine whether it is reliable.
@@ 3.2 Problem of Interest
Figure 1: Illustration of the detection of sensor attacks.
For each sensor, the evaluation of the observation sequence relies on the inference of each observation. Under Assumption 1, $y_{t}$ is considered normal when $p_{t}=0$, as the observation noise conforms to a Gaussian distribution; whereas when $p_{t}=1$, the observation noise is with unknown distribution, which is considered abnormal. Based on the number of occurrences of $p_{t}=1$ within the observation indicators $p_{0:N}$, we can classify the sequential observations into three status: in normal case, with general outliers, or under malicious sensor attacks.
@@
The problems of obtaining $\{\hat{p}_{t}\}$ and $\{\hat{x}_{t}\}$ are coupled. A reliable detector relies on accurate state estimator, while at the same time, effective elimination of anomalies contributes to better estimation performance. We formulate it as a dual-variable optimization problem.
@@
It is a Mixed-Integer Programming (MIP) problem with continuous variables ($\hat{x}_{0:N}$) and Boolean variables ($\hat{p}_{0:N}$). Solving such problem is challenging due to the non-convexity of the objective function and the combinatorial complexity introduced by integer variables.
@@
4 State Estimation with Secure Data: A Direct Computational Approach
In this section, we first solve a special case of Problem 1, i.e., when the observation indicators are fixed, what is the closed-form solution of equation (7)? Furthermore, when observation indicators changes, how should the solution of equation (7) be updated? These are the two key steps of our algorithm.
When $\hat{p}_{0:N}$ is fixed, we can denote the selected index set as $\mathcal{O}=\{i|\hat{p}_{i}=0\}$, and then Problem 1 is reduced into the following equivalent problem.
@@ 4 State Estimation with Secure Data: A Direct Computational Approach
As mentioned in Section 2.3, the solutions for MAP and LMMSE estimation in linear Gaussian systems are the same. Therefore, in this section, we present a direct solution to equation (8) from the perspective of LMMSE.
4.1 Sequential Estimation with Partial Observations
@@
Define the column vectors $X=col\{x_{0},\cdots,x_{N}\},Y=col\{y_{0},\cdots,y_{N}\}$ and $\hat{X}_{\mathcal{O}}=col\{\hat{x}_{0|\mathcal{O}},\cdots,\hat{x}_{N|\mathcal{O}}\}$. The following formulation describes the sequential estimation problem with partial observations.
@@
[…]
where each element in $X$ and $Y$ satisfies,
$x_{t+1}$ $\sim\hskip 0.0pt\mathcal{N}(Ax_{t},Q),\hskip 0.5ptt=0,\cdots,N,$
@@
This problem is equivalent to finding the optimal estimation for each state individually. Based on the linear estimation theory, the information from $y_{\mathcal{O}}$ can be well utilized by the Kalman filtering process and RTS smoother. It is worth mentioning that for those states missing corresponding observations, we can skip the step of measurement update.
@@
    Calculate $\{K_{i}\}$$\{F_{i}\}$ according to the Kalman filter and RTS smoother process. For those intermittent observations that $\{y_{j}|j\notin\mathcal{O}\}$, set the Kalman gain to zero $\{K_{j}=0|j\notin\mathcal{O}\}$.
@@
Moreover, Lemma 1 provides the explicit linear relationship between estimated states $\hat{x}_{0:N|\mathcal{O}}$ and the observations $y_{\mathcal{O}}$. For those $\{y_{i}|i\notin\mathcal{O}\}$, when multiplied by the matrix $K_{\mathcal{O}}$, the corresponding weights are zeros, which clearly indicates that they are unrelated to the estimate.
4.2 Estimation Update with New Observations